OpenStudy (anonymous):

Solve the system of equations. 4x + 7y = 6 and 6x + 5y = 20 someone please solve this. i have no idea how to do this

OpenStudy (anonymous):

which method will you like e.g. elimination, substitution ...

OpenStudy (anonymous):

elimination

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Same method you can apply.. Multiply the 1st equation by 6 and second equation by 4.. Subtract both the equations and you will have the value for y and then you can easily find the value of x...

OpenStudy (anonymous):

1. 4x + 7y = 6 *5 2. 6x + 5y = 20*7

OpenStudy (anonymous):

then by eliminating y we get 22x = 110

OpenStudy (anonymous):

x = 110/22

OpenStudy (anonymous):

x = 5

OpenStudy (anonymous):

then can you find the rest

OpenStudy (anonymous):

4x+7y=6 ,6x+5y=20 find the lcm of either the x terms or the y terms. the lcm of 4,6=12 3(4x+7y=6),2(6x+5y=20) u get 12x+21y=18,12x+10y=40 subtractin d equations v get 21y-10y=22,11y=-22,y=-2 x=5

OpenStudy (anonymous):

thank yall for the the help ! :)

OpenStudy (anonymous):

24x+42y=36 24x+20y=80 Then Subtract one and second equation: 22y=-44 And y = -44/22= -2 And therefore x = 5...

OpenStudy (anonymous):

x=5, y=-2

OpenStudy (anonymous):

:)D

OpenStudy (anonymous):

Please allow me to add: When you are solving for a system of two linear equations such as your system of equations, you are basically being asked if the two lines intersect and if they do where they intersect. So you could have also solved this system of equations graphically. You would have graphed each line and you would get the point of intersection at (5, -2) Recall: Your system would have had no solution if the lines were parallel (given they were not the same line) since two parallel lines do not intersect ex: y= 1/2 x and y= 1/2 x -1 NO SOLUTION and you would have had infinitely many solutions if your equations would have boiled down to being the same line 3y=3x-3 and y=x-1 INFINITELY MANY SOLUTIONS all the points on the line

OpenStudy (maheshmeghwal9):

First find the value of x. Than put this value in any equation. d^_^b